# Eye-tracking-Driven Shared Control for Robotic Arms: Wizard of Oz Studies to Assess Design Choices

### v1

@@ 2.4. Participation in assistive robot design
Aktion Mensch, a german organization for social projects, states that current assistive devices often fail to meet the needs of users (Borgstedt and Möller-Slawinski, 2023). In eye-tracking controls most studies recruit non-disabled people as participants (Fischer-Janzen et al., 2024). Reasons why people with disabilities are rarely included in studies are recruitment challenges, small sample sizes, and transportation logistics (Nanavati et al., 2024a). As shown by Stalljann et al., control inputs are experienced differently by disabled and non-disabled participants (Stalljann et al., 2020). Including people with disabilities helps in the design process of assistive devices by gaining insight into their lives and needs (Baumeister et al., 2021; Stanger et al., 1994; Nanavati et al., 2024a).
Nanavati et al. (Nanavati et al., 2024a) discusses approaches based on three-dimensional key features to find the most appropriate study design for the goal of the study. The first dimension decides between individual or community-level participation. With individuals, the chance of a deeper insight into the user’s life and preferences can be gained, while community work can lead to a broad insight from many participants. It also affects the time needed to conduct the studies and the connection between participant and experimenter. The second dimension highlights the logistical burden on participants and researchers. People with severe physical disabilities are usually dependent on their caregiver. Depending on the robot setup, it is easier to conduct studies at the participant’s location, to reduce the need of special transportation and availability of caregivers. The third dimension describes the benefits to researchers and the community. Participants have an intrinsic motivation to participate in the design process of assistive devices and robots. They can benefit by providing impetus to adapt the systems to their needs. Researchers benefit by potentially making important connections for future work.
Although it does affect the length of the study, it is worthwhile to involve people with disabilities in the design process. These findings influenced the study design in Section 3.4.
@@ 3. Methods
In the introduction and background, we presented influencing factors and user challenges that affect the adoption rates of robotic arm systems. These challenges and factors were partially incorporated into the design of the system, presented in Section 3.1. Several groups of interest were interviewed with two study designs, to obtain broader feedback on the shared control design. The setup described in Section 3.3 and 3.2 was used in both studies. An online survey was sent to stakeholder groups, as well as their families and medical professionals and is presented in Section 3.4. Measurements were conducted to verify technical functionality and the reaction of participants interacting with the robot in person (see Section 3.5).
@@ 3.1. Shared control design choices
The initial idea of the system is to include a shared control in which parts of the robot movement are automated, such as trajectory planning, grasping and object manipulation. The level of autonomy is presented following the guidelines of Beer at a. (2014) (Beer et al., 2014).
What task is the robot to perform? This shared control is designed to assist people with severe physical disabilities in everyday tasks. Ongoing work is realizing a framework that can be extended to include everyday activities previously unknown to the robot. The first approach includes the tasks of grasping and handing drinks, filling a glass with a drink, and grasping and placing objects in locations selected by the user. These tasks can be performed in unsorted environments such as at home and in public places, e.g. a restaurant.
@@
These tasks represent two of the most preferred tasks for robotic assistance in everyday life, as shown in Section 2.2.
@@
What aspects of the task should the robot perform? The chosen input modality is eye-tracking, since eye movement is often unrestricted or only slightly restricted in contrast to other motor functions such as arm movement in cases of locked-in syndrome or cerebral palsy. The control is realized by fixating the desired object and the robot decides based on the selected object which available task to perform. Multiple objects can be selected by gaze to make task selection more robust. For example, looking at a cup could indicate that the user wants to drink from it, move it, fill it, or combine available tasks. By adding the ability to combine the cup with other objects, such as selecting a bottle or a location, the task can be distinguished into refilling the cup or placing it in the selected location.
Based on object and ask information, the robot performs the task autonomously. It determines the location of the object, a suitable grasp related to the task, and calculates the optimal trajectory.
@@
To what extent can the robot perform those aspects of the task? The user is always in control of the robot. Emergency stop methods are considered to ensure the well-being of the user. Validation methods are planned to intervene in case of incorrect task execution before the robot starts the task. Settings are planned to reduce the robot’s velocity if desired. With the implementation of dwell time selection of objects, incorrect interaction with the object is reduced with natural gaze behavior, thus minimizing the Midas Touch problem.
@@
What level can the robot’s autonomy be categorized? According to the classification of Beer et al. (Beer et al., 2014) robot autonomy is categorized as Shared Control with Human Initiative.
@@
How might autonomy influence HRI variables? Reliability and functionality of the system must be given. Safety for the user must be ensured and is part of ongoing discussions. This system is designed to be easy to learn and use. Since there is no physical interface such as a display, the correct selection of the object depends on the accuracy of the eye tracker. Objects are selected by focusing on them in real-world settings. If the eye tracker is incorrectly calibrated, the gaze may not be tracked to the object due to high accuracy errors. This raises the question of how serious this error is in this setup, which will be tested with the Wizard of Oz setup. A wearable eye-tracker is used as input device. Influences on social variables such as aesthetics or the occupation of the voice by voice control, which could lead to a rejection of the system, are minimized.
@@ 3.2. Wizard of Oz design
Based on the results of the literature review, we decided to do an early evaluation of the design if it meets the expectations for the influencing factors given by the community. This approach has the advantage of shortening the development time by intervening in case of low acceptance or functional challenges. Therefore, the in the following used system represents a Wizard of Oz design, where the autonomously controlled robot tasks are performed by the experimenter, as seen in Figure 1. The objects were placed in a predefined position in the robot’s workspace. All trajectories needed to realize the automated task execution were hard-coded and could be selected by the experimenter clicking on the program. The safety of the participant was ensured by separating the workspace of the participant and robot.
The experimenter was informed about the task sequence of the user and observed the participant’s visualized eye movements in a live video on the monitor. As soon as the participant fixated on an object, the corresponding task was started via the robot’s web interface.
Five tasks were selectable with the Wizard of Oz system: Picking up food with a fork, turning on a light switch, picking up and placing a block, scooping from a bowl, and filling a glass with water and handing it to the participant. Each task was related to different objects, as shown in Figure 2. The selection were based on the initial idea for the system and the outcomes of the literature review.
A video of task performance was created to be implemented in the online survey to visualize the function of the system (Section 3.4). This video is included in the Supplementary Materials. The system was also used in the hands-on study, where participants operated the system in a controlled environment (Section 3.5). Both studies were reviewed and approved by the ethics committee of the first author’s institution.
@@
3.3. Technical setup
The Wizard of Oz setup consists of a robot (Kinova, Gen3 with a gripper Robotiq, 2F-85) controlled via the standard web interface provided by the manufacturer. Data were recorded using wearable eye-tracking glasses (Tobii, Tobii Pro Glasses 3, accuracy = 0.6). Gaze visualization, data analysis, and post-processing were performed using Tobii Pro Lab (Pro, 2014). The control computer to move the robot and visualize the participant’s gaze was placed 2 m behind the participant in the hands-on study so that the user was unaware of the experimenter’s role.
@@ 3.4. Online survey
The online survey was chosen as the data collection tool because it provided a standardized way to present information and collect feedback. The purpose is to gain broad feedback from people with disabilities, their family members and caregivers to assess the usability and accessibility of the system. The logistical challenges described by Nanavati et al. (Nanavati et al., 2024a) were minimized by providing an overview of the systems to the stakeholder group via online survey and video. It was advantageous because it allowed immobile people to participate through the use of smart devices. People living at home could participate, as opposed to a real-world demonstration where the authors could only visit a few facilities. In addition, it offered the convenience of completing the survey at the participant’s preferred time, accommodating those with complex schedules or who require assistance. Due to the anticipated limited availability of accessible computer controls, an extension was designed for family members and caregivers to obtain more comprehensive feedback. The purpose of this online survey was to answer the following questions:
@@
  • Which tasks are preferred so that the robot has the greatest benefit in everyday life?
@@ 3.4.1. Distribution method
The survey was available online from September 18, 2023 to January 22, 2024. 32 organizations in Germany, Switzerland, and Austria were contacted by email to reach the stakeholder groups. With reference to the distribution channels mentioned by Nanavati et al., we selected the following organizations and platforms (Nanavati et al., 2024a). Seven organizations shared the survey, such as FGQ (Paraplegic Support Association), LIS e.V. (Locked-In Syndrome Association), Diakonie (social services), Lebenshilfe (non-profit association), and the Swiss Paraplegic Centre. On social media, the survey was shared a total of three times by the authors’ private accounts and by their institutions. FGQ shared it in their private Facebook group and a post was made in the private forum of the Swiss Paraplegic Centre.
Figure 3. Distribution of questions asked for each group. Q outlines the questions’ context and the number of the questions for each category is given.
[…]
3.4.2. Survey structure
The online survey consisted of 31 questions divided into three groups and a video describing the eye-tracking control of the robotic arm structured as shown in Figure 3. The video showed the calibration process, the light switch task, the pick and place task, and the scoop task. In the survey, people with disabilities were asked questions to assess their level of disability and daily performance. Their need of the control modality, their access to assistive devices, and their evaluation of the presented eye-tracking control was surveyed. Open-ended questions were used to identify tasks that the robot should be able to perform. Family members and caregivers were asked about the amount of time they spend providing care and their estimation of the number of people with physical disabilities in their care facility. In contrast to the questions above, family members and caregivers were first asked what tasks they would like a robot to perform in general, regardless of the type of robot. In this situation, the participants were unaware of the robotic system. Then the system was presented and questions similar to the first questionnaire were asked. Finally, all groups were given the opportunity to provide feedback on the survey and the system itself.
All of the questions asked in the questionnaire, along with the participants’ responses, are available in the Supplementary Materials. The survey was designed with LimeSurvey.
@@ 3.4.3. Evaluation method
A frequency analysis was performed to evaluate the users’ perception of the system and the desired tasks. The tasks were categorized as introduced in Section 2.2: tasks related to eating, drinking, picking up objects, getting along, personal hygiene, and miscellaneous tasks.
@@ 3.5. Hands-on user study
In addition to the stakeholder evaluation of the system, the Wizard of Oz design was used to analyze real-world interaction. Evidence from the online survey suggests that participants are able to use the system based on their motor function and understanding of the system. The system realization might pose challenges in terms of functionality due to insufficient accuracy or natural gaze behavior. Participants were asked to interact with the robot to record gaze data that would be used to evaluate hardware constraints. Questions answered by this study include:
@@
  • What are the technical difficulties with the design?
  • Does participant perception vary greatly between the hands-on study and the online survey?
@@ 3.5.1. Procedure and participants
24 participants took part in the study (7 female/17 male, mean age = 30.5, 11 non-prior experience group/13 prior-experience group with robots) to minimize the influence of learning effects or behavioral adaptation related to the 24 emerging permutations in the task sequence. Each participant was informed and gave consent to be included. They were asked to sit in a chair in front of the robot workspace. The setup and calibration of the eye-tracker was performed as described in (Pro, 2024). Task start was selected by looking at one of two objects in the drinking, eating, and picking and placing tasks (see Figure 1). Participants chose the objects according to their preference. The experimenter performed the role of the automation as described in Section 3.2.

### v2

@@ 2.4. Participation in assistive robot design
The previous section presented characteristics that influence user acceptance and robot abandonment. Several authors argue that involving the community and people with disabilities in the design of assistive devices is more advantageous than involving people without disabilities. Including PSMD in the design process of assistive devices helps designers gain insight into the lives and needs (Baumeister et al., 2021; Stanger et al., 1994; Nanavati et al., 2024a). Arboleda et al. point out that end users’ perspectives can differ from those of other groups, such as caregivers (Arevalo Arboleda et al., 2021). As Stalljann et al. demonstrate, control inputs are experienced differently by disabled and non-disabled participants (Stalljann et al., 2020). Most studies on eye-tracking controls recruit non-disabled people as participants (Fischer-Janzen et al., 2024). Reasons PSMD are rarely included in studies include recruitment challenges, small sample sizes, and transportation logistics (Nanavati et al., 2024a).
Nanavati et al. (Nanavati et al., 2024a) discuss approaches based on three-dimensional key features to determine suitable study designs for achieving the study’s objective. The first dimension determines whether participation is at the individual or community level. Individual participation provides the opportunity for deeper insight into the user’s life and preferences, while community work provides insight from many participants. This dimension also affects the time required to conduct the study and the connection between participants and experimenters. The second dimension highlights the logistical effort required of participants and researchers. PSMD participants may require assistance with transportation to study sites, which can facilitate conducting studies at the participants’ locations. Arboleda et al. stated that conducting the study itself might disrupt participants’ daily routines, which could induce stress (Arevalo Arboleda et al., 2021). Therefore, other possibilities can be considered, such as online interviews. Depending on the robot setup, it may be easier to conduct studies at the participant’s location to reduce the need for special transportation and availability of medical professionals. The third dimension describes the benefits to researchers and the community. PSMD, their families, and medical professionals are intrinsically motivated to participate in the design process of assistive devices and robots. They can benefit and contribute by providing insight to adapt the systems to their needs.
These findings demonstrate the value of involving PSMD in the design process and influenced the study design in Section 3.3.
@@ 3. Methods
In Sections 1 and 2, we presented the influencing factors and user challenges that affect acceptance of eye-tracking-driven shared controls. These factors and challenges were incorporated into the system design presented in Section 3.1. Several groups of interest were interviewed with two study designs, to obtain broader feedback. The setup described in Section 3.2 was used in both studies. An online survey was sent to stakeholder groups, their families and medical professionals. The results are presented in Section 3.3. Measurements were conducted to verify technical functionality and participants’ perception of the system interacting directly with the robot (see Section 3.4).
@@ 3.1. Shared control design choices
The initial idea for the system is to include a shared control, which automates parts of the robot movement, such as trajectory planning, grasping, and object manipulation. The level of autonomy is presented according to the guidelines of Beer et al. (2014) (Beer et al., 2014). The following considerations were used to create the Wizard of Oz design and describe the interaction to the participants.
What task is the robot to perform? This shared control system is designed to assist PSMD with daily tasks. Ongoing work is developing a framework that can be expanded to include new ADL, thereby increasing the range of tasks it can perform. The first approach includes the tasks of picking up food with a fork, turning on a light switch, picking up and placing a block, scooping from a bowl, pouring water into a glass, and handing it to the participant. These tasks have to be performable in unstructured environments, such as at home and in public places like restaurants.
@@
As shown in Section 2.3, these tasks are preferred for robotic assistance in everyday life.
@@
What aspects of the task should the robot perform? The chosen input modality is eye-tracking, because eye movement is often unrestricted or only slightly restricted in cases of locked-in syndrome or cerebral palsy, in contrast to other motor functions, such as arm movement. The robot is controlled by fixating the desired object, and the robot decides which available task to perform based on the selected object. Selecting multiple objects with the gaze makes task selection more robust. For instance, looking at a cup could indicate that the user wants to drink from it, move it, fill it, or combine available tasks. Adding the ability to combine the cup with other objects, such as a bottle or a location, distinguishes the task as refilling the cup or placing it in the selected location. Based on the information about the object and the task, the robot performs the task autonomously. The robot determines the object’s location, selects a suitable grasp related to the task, and calculates the optimal trajectory.
@@
To what extent can the robot perform those aspects of the task? The user is always in control of the robot in order to maintain a sense of agency (discussed in (Nanavati et al., 2024b)). Validation methods are planned to intervene in case of incorrect task execution before the robot begins (suggested in (Baptista et al., 2025)). Settings are planned to reduce the robot’s velocity, if desired (suggested by (Bhattacharjee et al., 2020)). Implementing dwell time for selecting objects has the potential to reduce incorrect interactions with objects through natural gaze behavior, thus minimizing the Midas Touch problem.
@@
What level can the robot’s autonomy be categorized? According to the classification of Beer et al. (Beer et al., 2014), robot autonomy is categorized as Shared Control with Human Initiative.
@@
How might autonomy influence HRI variables? The system must be reliable and functional. User safety is part of ongoing discussions and must be ensured. The system is designed to be easy to learn and use. In real-world settings, objects are selected by focusing on them. However, the accuracy of the eye tracker limits the correct selection of the desired object when gaze is tracked outside the area of interest (AoI). This raises the question of how accuracy errors influence the functionality of this approach. This will be tested using the Wizard of Oz setup.
@@ 3.2. Wizard of Oz design
An early system evaluation was conducted to determine whether the design meets the community’s expectations for the found influencing factors. This approach shortens development time by intervening early in the event of low acceptance or functional challenges. In the Wizard of Oz design, the experimenter performs the autonomously controlled robot tasks, as seen in Figure 1. Table 1 presents the Wizard of Oz system according to the guidelines presented by Riek (Riek, 2012). Figure 2 shows the five selectable tasks with the Wizard of Oz system. A video of task performance was created and implemented in the online survey to demonstrate the system’s functionality (Section 3.3). This video is included in the Supplementary Materials. The system was also used in the hands-on study, where participants operated the system in a controlled environment (Section 3.4). Both studies were reviewed and approved by the ethics committee of the first author’s institution.
Table 1. Wizard of Oz presentation after Riek (Riek, 2012) stating experimental components to the robot, user, wizard and to general information.
Experimental Component Online survey Hands-on study
Robot 7-DoF Kinova Gen3
Level of autonomy Teleoperation
Robot capabilities Food item, scooping food, filling cup Food item, pick and place, filling cup, light switch
No. of participants 34 24
Demographics Age: $40\pm 27$; Age: $30.5\pm 8$;
14 PSMD, 9 family members, 11 medical professionals; 11 non-prior experience group, 13 prior-experience group;
14 male, 17 female, 3 diverse 7 female,17 male
Instructions See Section 3.3.3 See Section 3.4.1
Convincing simulation Not asked Yes
Wizard Experimenter took the role of the wizard
Familiar with hypothesis? Yes
Production variables - See Section 3.4.1
Recognition variables - See Section 3.4.1
Wizard error Video representation, no variation between tests Experiments were repeated when robot malfunctioned, occurrence: 1.
Location Unknown, not asked due to privacy regulations At the laboratory
Environmental constants Unknown Setup as shown in Figure 2
Scenario As shown in video (see Supplementary Materials) See additional information to scenario.
Iterative design process No No
Limitations See Section 6
Reason for omitting the item ”Experimental Component” from the Wizard presentation of Riek et al.: The hypothesis for the robot was to reproduce each participant’s trajectory and create as little variation as possible. Thus, the robot’s software computed the trajectory planning. The points between the trajectories were programmed so that the overall task trajectory would be reproducible between participants. As Arboleda et al. indicated, robot behavior is perceived differently in virtual and real presentations (Arevalo Arboleda et al., 2021). Therefore, a crucial aspect of the behavioral hypothesis was to compare the two presentations. Distinct reactions of the participants were not previously hypothesized, since the goal of this work was to gain insights into user expectations and behavior.
Additional information on the item ”Scenario”: The objects were placed in predefined positions within the robot’s workspace. All of the trajectories necessary for the automated task execution were hard-coded and could be selected by the experimenter by clicking on the program. The participant’s safety was ensured by separating their workspace from the robot’s.
@@ 3.3. Online survey
The online survey was chosen as the data collection tool because it provided a standardized way to present information and collect feedback. The goal was to obtain comprehensive feedback from PSMD, their family members, and medical professionals in order to evaluate the system’s usability and accessibility. Providing an overview of the system to the stakeholder group via online survey and video minimized the logistical challenges described in Section 2.4. This allowed immobile individuals to participate using smart devices. Unlike a real-world demonstration, where the authors could only visit a few facilities, people living at home could participate. Additionally, the survey could be completed at the participant’s preferred time, which was convenient for those with complex schedules or who required assistance. Due to the anticipated limited availability of accessible computer controls, an extension was designed for family members and medical professionals to obtain more comprehensive feedback. The purpose of the online survey was to answer the following questions:
@@
  • Which tasks should be prioritized to maximize the robot’s usefulness in everyday life?
@@
3.3.1. Survey development and structure
The survey focused on investigating five topics, which are listed in Table 2. First, questions were asked about the participants’ demographics to provide comparable and reusable baseline data. After the five topics were covered, participants were given the opportunity to provide feedback on the presented system and make general remarks. As stated in Section 2.4, perceptions of assistive devices can differ between end users and individuals in close contact with them, such as medical professionals and family members. Our goal was to receive a broad perspective on the prototype, which is why the survey varied between these three groups. Questions about the support situation were designed to evaluate the survey’s accessibility. Poor accessibility can be indicated by how Q4(PSMD) is answered. A higher number of relatives and medical professionals who did not involve the person they care for can indicate time constraints. Furthermore, if everyone in Q4(relatives/friends) had stated that the person they care for had answered the survey, the answers would need to be weighted due to potential bias.
The topics of disabilities and assistive devices focus on whether the respondent is the intended audience. We implemented questions to reduce responses from people who misidentified their group affiliation. The topics also focus on whether the respondent is familiar with assistive devices, including robotic arms. The literature contains cases in which users reacted differently to devices when they had prior knowledge. This may affect their reaction to the presented system. The questions regarding the robot presentation were developed to identify signs of abandonment. Robotic arms are not widespread. Upon inquiry, two companies that sell assistive robotic arms approved as medical devices reported selling around 1,200 models in Germany, Switzerland, and Austria. Since this strongly affects the system’s perception, participants were asked if they had experience with robotic arms.
Lastly, task preferences were requested. As technology continues to evolve, the aim was to update the ongoing literature. As presented in Section 2.3, task preferences change quickly with the advent of new technology, such as smart home applications and hands-free smartphones. Some tasks, like using a fax machine, are no longer relevant. We asked participants to rank the importance of three tasks against each other.
Throughout the survey, we chose open-ended questions to receive rationale for the participants’ answers to yes/no questions, such as reasoning for being unable to use the system. After initial development, the survey was discussed and refined.
Table 2. Breakdown of topics asked in the survey and related questions.
Question topic PSMD Rel./Fr. Med. prof.
Support situation Q4+Q5 Q4+Q5 Q4
Disability & assistive devices Q6-Q8 Q6+Q7 -
Robotic arm use Q9-Q12 Q9+Q10 Q9+Q10
WoZ satisfaction Q13+Q14 Q11+Q12+Q14+Q15 Q11+Q12+Q14+Q15
Task preferences Q15 Q13 Q5+Q13
As a result, the online survey consisted of 31 questions, divided into three groups, as well as a video describing the eye-tracking control of the robotic arm, structured as shown in Figure 3. In addition to the tasks, the video showed the calibration process. The exact questions asked are provided in Appendix A. All survey questions, along with the participants’ responses, are available in the Supplementary Materials. The survey was designed using LimeSurvey.
Figure 3. Distribution of questions asked for each group. Q outlines the context of the questions and the number of the questions for each category is given.
@@ 3.3.2. Distribution method
The survey was available online from September 18, 2023, to January 22, 2024. 32 organizations in Germany, Switzerland, and Austria were emailed to reach the stakeholder groups. Seven organizations, including FGQ (Paraplegic Support Association), LIS e.V. (Locked-In Syndrome Association), Diakonie (social services), Lebenshilfe (nonprofit association), and the Swiss Paraplegic Centre, shared the survey. The authors’ private accounts and their institutions shared the survey a total of three times on social media. FGQ shared it in their private Facebook group, and the Swiss Paraplegic Centre posted about it in their private forum.
3.3.3. Participant instructions
First, participants were informed about the content and purpose of the study. Then, they were asked to give their consent to participate. Participants were then guided through the survey. Participants could skip questions at any time if they felt uncomfortable answering them. The video played when clicked on, and could be repeated if necessary. To offer different resolutions, participants could also open the video on YouTube.
@@ 3.3.4. Evaluation method
A frequency analysis was performed to evaluate users’ perceptions of the system and their desired tasks. As introduced in Section 2.3, the tasks were categorized as follows: eating, drinking, picking up objects, environmental interactions, personal hygiene, and miscellaneous.
@@ 3.4. Hands-on user study
In addition to the stakeholder evaluation of the system, a Wizard of Oz design was employed to analyze real-world interactions. Evidence from the online survey suggests that participants can use the system based on their answers. However, challenges affecting the system’s functionality may occur due to the eye tracker’s insufficient accuracy or natural gaze behavior. Participants interacted with the robot to record gaze data, which was then used to evaluate hardware constraints. This study addresses the following questions:
@@
  • What technical difficulties arise from the design?
  • Does perception vary between the hands-on study and the online survey?
@@ 3.4.1. Procedure and participants
To minimize the influence of learning effects or behavioral adaptation, 24 participants were involved, creating 24 emerging permutations in the task sequence. Each participant provided informed consent after the procedure was explained. They were asked to sit in a chair in front of the robot workspace. The eye-tracker was set up and calibrated as described in (Pro, 2024). Participants started the task by fixating on one of two objects in the drinking, eating, light-switching, and picking-and-placing tasks with gaze (see Figure 1). Participants chose the objects according to their preference. The experimenter performed the role of the automation, as described in Section 3.2.
Production and recognition variables were discussed prior to the test. Within the first second of selection, the wizard selected the corresponding task via the web interface. The robot’s reaction took between one and five seconds from the time the first reaction was recognized, including the wizard’s reaction time. As shown in Table 1, one wizard error occurred, causing the robot to stop shortly after the experiment started due to an incorrect mode selection. This experiment’s results were discarded, and it was repeated with another participant to ensure naivety to the procedure and function of the Wizard of Oz setup.
Wizard training: The experimenter was informed of the user’s task sequence and observed the participant’s eye movements in real time on the monitor. As soon as the participant fixated on an object, the corresponding task started via the robot’s web interface.